OpenStudy (anonymous):

find a Differential equation for the family of this curve... cy^2+4y=2x^2 c=constant

OpenStudy (anonymous):

see if you can use implicit differentiation on it and solve for y'

OpenStudy (anonymous):

what the procedure last time I diferentieted implicitly was last year

OpenStudy (anonymous):

f(y) -> f'(y) * y' eg. d/dx (4y^3) = 12y^2 *y'

OpenStudy (anonymous):

f(x) -> f'(x) (as usual) eg. d/dx (x^3) = 3x^2

OpenStudy (anonymous):

ow jah I tried dat but I do not get the given answer..

OpenStudy (anonymous):

what was the final answer?

OpenStudy (anonymous):

it's suppose to be(x^2-y)y'=xy

OpenStudy (anonymous):

hmh, let me look it over a bit...

OpenStudy (anonymous):

ah, I think we have to use partials..

OpenStudy (anonymous):

\[2cyy^'+4y^'=4x\] \[y ^{'}=\frac{ 2x }{ cy+2}\]

OpenStudy (anonymous):

jonask can u change yr answer to the 1 I gave?

OpenStudy (anonymous):

coz u dd exactly wat i dd....jst tht the answr do not dorrespond!

OpenStudy (anonymous):

I think I messed up @ekim2012 , but I found the right procedure and I'm looking it over. Just a moment..

OpenStudy (anonymous):

cul

OpenStudy (anonymous):

i divided by 2 \[cyy^'+2y^'=2x\] factorise then \[y^'(cy+2)=2x\]

OpenStudy (anonymous):

no..not coreect answer!!!!:)

OpenStudy (anonymous):

you sure you're on the right solution @ekim2012 ? I did it the other way and got the same result as using implicit diff.

OpenStudy (anonymous):

yes..but it's fine unkess they wrote the wrong solution..but I dn't thnk so!!

OpenStudy (anonymous):

Are you sure you gave the right eqn. ?

OpenStudy (anonymous):

you can also solve for x\[x^2=cy^2+4y\] \[x=(cy^2+4y)^{1/2}\] \[\frac{ dx }{ dy }=\frac{ 1/2(2cy+4) }{ \sqrt{cy^2+4y} }=\frac{ cy+2 }{\sqrt{cy^2+4} }\]

OpenStudy (anonymous):

4y inside,to find dy/dx you can invert

OpenStudy (anonymous):

is the solution only in terms of x

OpenStudy (anonymous):

k..I'll just ask my lecturer tommorow nd report to u guys...yes it's only in terms of x...here is it (x^2-y)y'=xy

OpenStudy (anonymous):

Sorry, @ekim2012 ! It's possible that I'm thinking about this incorrectly...

OpenStudy (anonymous):

fine....

OpenStudy (experimentx):



OpenStudy (anonymous):

Thanks experiment....u gud